OpenStudy (anonymous):

4x^2-3y^2=-8 5x^2+2y^2=13 solve by the addition method

OpenStudy (anonymous):

\[8 x^2-6 y^2=-16\]\[15 x^2 + 6 y^2 = 39 \]--------------- 23 x^2 = 23